# The twisted coarse Baum–Connes conjecture and relative hyperbolic groups

### v1

@@ Abstract
In this paper, we introduce a notion of stable coarse algebras for metric spaces with bounded geometry, and formulate the twisted coarse Baum–Connes conjecture with respect to stable coarse algebras. We prove permanence properties of this conjecture under coarse equivalences, unions and subspaces.
@@ 1 Introduction
is an isomorphism, where the left side is the coarse $K$-homology and the right side is the $K$-theory of the Roe algebra of $X$. This conjecture has many important applications in geometry and topology, including the Gromov–Lawson conjecture on the existence of Riemannian metrics with positive scalar curvature, and the Novikov conjecture on the homotopy invariance of higher signature on oriented closed manifolds. It has been verified for a large class of metric spaces, for example, metric spaces which admit a coarse embedding into Hilbert space [Yu00], and some metric spaces which are not coarsely embeddable into Hilbert space, including relative expanders [DWY23].
To study the coarse Baum–Connes conjecture for the extension $1\to N\to G\to G/N\to 1$ with $N$ and $G/N$ coarsely embeddable into Hilbert space (cf. [AT19]), the first author and L. Guo [DG24] developed a twisted coarse Baum-Connes conjecture with coefficients. In their work, they introduced the notion of a coarse algebra. Roughly speaking, for a metric space $X$ with bounded geometry, a $C^{*}$-algebra $\mathcal{A}$ and the algebra of all compact operators $\mathcal{K}$ on a separable infinite dimensional Hilbert space $\mathcal{H}$, a coarse algebra $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$ is a $C^{*}$-subalgebra of $\ell^{\infty}(X,\mathcal{A}\otimes\mathcal{K})$ which is invariant under the pullback by partial translations. This is a coarse geometric counter-part of a coefficient of a crossed product $C^{*}$-algebra. Then they formulated the twisted Roe algebra $C^{*}(X;\Gamma(X,\mathcal{A}\otimes\mathcal{K}))$ associated with a coarse algebra $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$. Using this framework, they verified the coarse Baum-Connes conjecture holds for a metric space with a coarse fibration for which each fiber and the base space are coarsely embeddable into Hilbert space, and this applies to the extension of coarse embeddable groups.
Based on the aforementioned framework of the twisted coarse Baum–Connes conjecture, we introduce a notion of stable coarse algebra for a metric space with bounded geometry and define a refined version of twisted coarse Baum–Connes conjecture. This refined version of twisted coarse Baum–Connes conjecture has better permanence properties, for instance, inheritance to subspaces and coarse invariance. We say that a coarse algebra $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$ is stable if it is invariant under the multiplication in $B(\mathcal{H})$-direction and under forming a matrix (Definition 2.1). We can generalize the twisted Roe algebra to non-discrete metric spaces with bounded geometry. Let $M$ be a metric space and $X\subset M$ a net in $M$. Given a stable coarse $X$-algebra $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$, we can define the $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$-twisted Roe algebra $C^{*}(M;\Gamma(X,\mathcal{A}\otimes\mathcal{K}))$.
@@
The twisted coarse Baum–Connes conjecture [DG24] claims that the twisted coarse assembly map is an isomorphism with respect to any coarse $X$-algebra $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$. We study when it is isomorphism for any stable coarse algebras.
@@ 2 The twisted coarse Baum–Connes conjecture with respect to stable coarse algebras
In this section, we review the background on the coarse Baum–Connes conjecture together with its variants. More precisely, we will focus on a version of twisted coarse Baum-Connes conjecture with respect to certain stable coarse algebras. This is a new framework to study the coarse geometry of metric spaces with bounded geometry. Compared with the usual coarse Baum-Connes conjecture, the new framework captures more information of the large scale geometry of the considered space.
@@ 2.1 Twisted Roe algebras
    For each $x\in X$, we have that
    $\overline{\{f(x):f\in\Gamma(X,\mathcal{A}\otimes\mathcal{K})\}}=\mathcal{A}\otimes\mathcal{K};$
@@
    For any two bounded sequences $a,b\in\ell^{\infty}(X,B(\mathcal{H}))$, and $f\in\ell^{\infty}(X;\mathcal{A}\otimes\mathcal{K})$, we define a new function $f^{\prime}\in\ell^{\infty}(X;\mathcal{A}\otimes\mathcal{K})$ by
    $f^{\prime}(x)=\left(\text{id}_{\mathcal{A}}\otimes a(x)\right)f(x)\left(\text{id}_{\mathcal{A}}\otimes b(x)\right).$
    Then $f\in\Gamma(X,\mathcal{A}\otimes\mathcal{K})$ implies $f^{\prime}\in\Gamma(X,\mathcal{A}\otimes\mathcal{K})$ for any two bounded sequences $a,b\in\ell^{\infty}(X,B(\mathcal{H}))$;
@@
    The propagation of $T$ is defined to be
    ${\rm Propagation}(T):=\sup\{d(x,y):T_{x,y}\neq 0\}\in[0,\infty].$
    We say that $T$ has finite propagation if ${\rm Propagation}(T)<\infty$.
@@
Next, we recall the definition of twisted Roe algebras. This notion was first introduced by the first-named author together with L. Guo in [DG24], where the coefficient algebras were taken to be coarse algebras. In the present paper, we instead use stable coarse algebras to define twisted Roe algebras. With this modification, the resulting twisted Roe algebras are invariant under coarse equivalences in a certain sense.
We fix a net $X$ of $M$ and a Borel covering $\{B_{x}\}_{x\in X}$ satisfying the four conditions in Lemma 2.4. For every $T=\left(T_{xy}\right)_{x,y\in M_{0}}\in B(\ell^{2}(M_{0},\mathcal{H}_{\mathcal{A}}))$ and every partial translation $v:D\to R$ on $X$, we define a bounded function
@@
in the following way. For every $x\in X$, fix a unitary
$U_{x}:\ell^{2}(B_{x}\cap M_{0})\otimes\mathcal{H}\to\mathcal{H}$ (1)
between Hilbert spaces. This induces an isomorphism on the Hilbert $A$-modules $U_{x}\otimes\text{id}_{\mathcal{A}}:\ell^{2}(B_{x}\cap M_{0})\otimes\mathcal{H}_{\mathcal{A}}\to\mathcal{H}_{\mathcal{A}}$. We define
@@
    For the stable coarse $X$-algebra $\ell^{\infty}(X,\mathcal{A}\otimes\mathcal{K})$, we have that
    $C^{*}(X,\ell^{\infty}(X,\mathcal{A}\otimes\mathcal{K}))=C^{*}(X,\mathcal{A}).$
@@
    For the stable coarse $X$-algebra $C_{0}(X,\mathcal{A}\otimes\mathcal{K})$, we have that
    $C^{*}(X,C_{0}(X,\mathcal{K}))\cong\mathcal{A}\otimes\mathcal{K}.$
@@
$f^{*}(\Gamma(Y,\mathcal{A}\otimes\mathcal{K}))=\{\phi\circ f:\varphi\in\Gamma(Y,\mathcal{A}\otimes\mathcal{K})\}$
Then, the proof of coarse invariance is same as that of Roe algebras, and in this case, it is not necessary to require that the coarse algebras are stable. In Theorem 2.14, we prove a more general coarse invariance.
@@
It is obvious that $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$ is a $C^{*}$-subalgebra of $\ell^{\infty}(X,\mathcal{A}\otimes\mathcal{K})$. For any partial translation $v$ on $X$ and any $\phi\in\Gamma(X,\mathcal{A}\otimes\mathcal{K})$, we show that $v^{*}\phi\in\Gamma(X,\mathcal{A}\otimes\mathcal{K})$. For any subset $\tilde{Y}\subset f(X)$, we fix a set of representatives
@@
$(v^{*}\phi)_{\tilde{X}}(y)=(v^{*}\phi)(x_{y})=\phi(v(x_{y}))$
and $(v^{*}\phi)_{\tilde{X}}(y)=0$ if $y\notin\tilde{Y}$ or $x_{y}\notin\operatorname{Dom}(v)$.
@@
$v_{j}^{*}(\phi_{\tilde{X}_{j}})(y)=\phi_{\tilde{X}_{j}}(v_{j}(y))=\phi_{\tilde{X}_{j}}(f(v(x_{y})))=\phi(v(x_{y}))$
and $v_{j}^{*}(\phi_{\tilde{X}_{j}})(y)=0$ if $y\notin A_{j}$. Therefore $v^{*}\phi=\sum_{j=1}^{N}v_{j}^{*}(\phi_{\tilde{X}_{j}})\in\Gamma(X,\mathcal{A}\otimes\mathcal{K})$. It is easy to see $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$ is stable if $\Gamma(Y,\mathcal{A}\otimes\mathcal{K})$ is. ∎
@@
Let $M$ be a metric space with bounded geometry, and let $X$ and $Y$ be nets of $M$ with associated Borel covers $\{B_{x}\}_{x\in X}$ and $\{C_{y}\}_{y\in Y}$ of $M$ satisfying the four conditions in Lemma 2.4. Suppose that $\Gamma(Y,\mathcal{A}\otimes\mathcal{K})$ is a stable coarse $Y$-algebra. If $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$ is a coarse $X$-algebra induced by any coarse equivalence between $X$ and $Y$, which is close to the identity map on $M$, by the formula (4), then we have an equality
@@
Let $M$ and $N$ be metric spaces with bounded geometry, and let $X\subset M$ and $Y\subset N$ be nets of $X$ and $Y$, respectively. Let $f:M\to N$ be a coarse equivalence, and fix Borel covers $\{B_{x}\}_{x\in X}$ and $\{C_{y}\}_{y\in Y}$ satisfying four conditions in Lemma 2.4, together with a coarse equivalence $g:X\to Y$ that is close to $f|_{X}$.
@@ 2.2 The twisted coarse Baum–Connes conjecture
Each Rips complex $P_{s}(X)$ can be equipped with a spherical metric as follows. The simplex $\left\{\sum_{i=0}^{n}a_{i}x_{i}\right\}$ spanned by $\{x_{0},x_{1}\dots,x_{n}\}$ can be identified with the positive quadrant $\left\{(c_{0},c_{1},\dots,c_{n})\right\}$ equipped with the standard Riemannian metric via the map
$\sum_{i=1}^{n}a_{i}x_{i}\mapsto\left(\frac{a_{0}}{\sum_{i=0}^{n}a_{i}^{2}},\frac{a_{1}}{\sum_{i=0}^{n}a_{i}^{2}},\dots,\frac{a_{n}}{\sum_{i=0}^{n}a_{i}^{2}}\right).$
Next, the Rips complex $P_{s}(X)$ is equipped with the path metric, and the distance between different components is defined to be infinity. For each $s>0$, we consider a countable dense subset $P_{r}(X)_{0}\subset P_{r}(X)$ such that $P_{s}(X)_{0}\subset P_{s^{\prime}}(X)_{0}$ if $s<s^{\prime}$.
@@
Since the right hand side of (6) is independent of $s$, passing to the limit yields the twisted coarse assembly map
@@
First, we discuss how a continuous and coarsely equivalent map $f:M\to N$ induces a map on the $K$-theory of their twisted localization algebras and show the coarse invariance of the twisted coarse Baum–Connes conjecture. We need the following terminology.
@@
Let $M$ and $N$ be metric spaces with bounded geometry. Fix their net $X\subset M$ and $Y\subset N$ and Borel covers $\{B_{x}\}_{x\in X}$ and $\{C_{y}\}_{y\in Y}$ of $M$ and $N$ satisfying the four conditions in Lemma 2.4, respectively. Let $f:M\to N$ be a continuous coarsely equivariant map. Fix any stable coarse $Y$-algebra $\Gamma(Y,\mathcal{A}\otimes\mathcal{K})$ and $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$ is the coarse $X$-algebra induced by $f$. Note that we can also take nets $X^{\prime}\subset M$ and $Y^{\prime}\subset N$ and Borel covers $\{B^{\prime}_{x}\}_{x\in X^{\prime}}$ and $\{C^{\prime}_{y}\}_{y\in Y^{\prime}}$ of $M$ and $N$, respectively such that $f$ is a bijection between $X^{\prime}$ and $Y^{\prime}$ and $f(B^{\prime}_{x})\subset C^{\prime}_{f(x)}$. We fix a decreasing sequence $\{\varepsilon_{j}\}_{j}$ of positive numbers with $\lim_{j}\varepsilon_{j}=0$. For each $x\in X^{\prime}$, since $f$ is continuous, by decomposing $C^{\prime}_{f(x)}$ into a smaller disjoint Borel sets satisfying the condition in Lemma 2.4, we obtain a unitary
@@
such that $\operatorname{supp}V_{j}(x)\subset\{(p,q)\in B^{\prime}_{x}\times C^{\prime}_{f(x)}:d(f(p),q)\leq\varepsilon_{j}\}$. We denote
@@
then $V_{j}$ is an $\varepsilon_{j}$-cover of $f$. We denote by $\Gamma(X^{\prime},\mathcal{A}\otimes\mathcal{K})$ and $\Gamma(Y^{\prime},\mathcal{A}\otimes\mathcal{K})$ the coarse algebra induced by the coarse equivalence between $X$ and $X^{\prime}$, and between $Y$ and $Y^{\prime}$, respectively. Then it is easy to see that the adjoint by $V_{j}$ sends $C^{*}(M;\Gamma(X^{\prime},\mathcal{A}\otimes\mathcal{K}))$ to $C^{*}(N;\Gamma(Y^{\prime},\mathcal{A}\otimes\mathcal{K}))$. But by Lemma 2.13, for originally fixed $\{B_{x}\}_{x\in X}$ and $\{C_{y}\}_{y\in Y}$, the adjoint by $V_{j}$ sends $C^{*}(M;\Gamma(X,\mathcal{A}\otimes\mathcal{K}))$ to $C^{*}(N;\Gamma(Y,\mathcal{A}\otimes\mathcal{K}))$. Define a uniformly continuous family of unitaries $(W(t))_{t\in[1,\infty)}$, where
@@
Then $(W(t))_{t}$ is a continuous cover of $f$ and the adjoint by $(W(t))_{t}$ intertwines $C^{*}_{L}(M;\Gamma(X,\mathcal{A}\otimes\mathcal{K}))$ and $C^{*}_{L}(N;\Gamma(Y,\mathcal{A}\otimes\mathcal{K}))$. We denote
@@
$\text{ad}(W(t))\left(C^{*}_{L}(M;\Gamma(X,\mathcal{A}\otimes\mathcal{K}))\right)\subset C^{*}_{L}(N;\Gamma(Y,\mathcal{A}\otimes\mathcal{K})).$

### v2

@@ Abstract
In this paper, we introduce a notion of stable coarse algebras for metric spaces with bounded geometry, and formulate the twisted coarse Baum–Connes conjecture with respect to stable coarse algebras. We establish permanence properties of the conjecture under coarse equivalences, as well as for unions and subspaces.
@@ 1 Introduction
is an isomorphism, where the left-hand side is the coarse $K$-homology and the right-hand side is the $K$-theory of the Roe algebra of $X$. This conjecture has many important applications in geometry and topology, including the Gromov–Lawson conjecture on the existence of Riemannian metrics with positive scalar curvature, and the Novikov conjecture on the homotopy invariance of higher signature on oriented closed manifolds. It has been verified for a large class of metric spaces, for example, metric spaces which admit a coarse embedding into Hilbert space [YU00], and some metric spaces which are not coarsely embeddable into Hilbert space, including relative expanders [DWY23].
To study the coarse Baum–Connes conjecture for the extension $1\to N\to G\to G/N\to 1$ with $N$ and $G/N$ coarsely embeddable into Hilbert space (cf. [AT19]), the first author and L. Guo [DG24] developed a twisted coarse Baum-Connes conjecture with coefficients. In their work, they introduced the notion of a coarse algebra. Roughly speaking, for a metric space $X$ with bounded geometry, a $C^{*}$-algebra $\mathcal{A}$ and the algebra of all compact operators $\mathcal{K}$ on a separable infinite dimensional Hilbert space $\mathcal{H}$, a coarse algebra $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$ is a $C^{*}$-subalgebra of $\ell^{\infty}(X,\mathcal{A}\otimes\mathcal{K})$ which is invariant under pullbacks by partial translations. This is a coarse geometric counterpart of a coefficient of a crossed product $C^{*}$-algebra. Then they formulated the twisted Roe algebra $C^{*}(X;\Gamma(X,\mathcal{A}\otimes\mathcal{K}))$ associated with a coarse algebra $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$. Using this framework, they verified that the coarse Baum-Connes conjecture holds for a metric space with a coarse fibration for which each fiber and the base space are coarsely embeddable into Hilbert space, and this applies to extensions of coarse embeddable groups.
Based on the aforementioned framework of the twisted coarse Baum–Connes conjecture, we introduce the notion of stable coarse algebra for a metric space with bounded geometry and define a refined version of twisted coarse Baum–Connes conjecture. This refined version of twisted coarse Baum–Connes conjecture has better permanence properties, such as inheritance to subspaces and invariance under coarse equivalence. We say that a coarse algebra $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$ is stable if it is invariant under multiplication in the $B(\mathcal{H})$-direction and under forming a matrix (Definition 2.1). We can generalize the twisted Roe algebra to non-discrete metric spaces with bounded geometry. Let $M$ be a metric space and $X\subset M$ a net in $M$. Given a stable coarse $X$-algebra $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$, we can define the $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$-twisted Roe algebra $C^{*}(M;\Gamma(X,\mathcal{A}\otimes\mathcal{K}))$.
@@
The twisted coarse Baum–Connes conjecture [DG24] claims that the twisted coarse assembly map is an isomorphism with respect to any coarse $X$-algebra $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$. We study when it is an isomorphism for any stable coarse algebras.
@@
The twisted coarse Baum–Connes conjecture with respect to any stable coarse algebras implies the usual coarse Baum–Connes conjecture. As an application, we can verify the coarse Baum–Connes conjecture for a larger class of metric spaces:
Corollary 1.4.
Let $G$ be a finitely generated group which is hyperbolic relative to a finite collection of subgroups $H_{1},H_{2},\dots,H_{n}$}. If each $H_{i}$ is an extension of groups admitting a coarse embedding into Hilbert space, then $G$ satisfies the coarse Baum–Connes conjecture.
Notably, the subgroups $H_{i}$ need not admit a coarse embedding into Hilbert space. This occurs, for example, in extensions constructed by G. Arzhantseva and R. Tessera in [AT19]. Such spaces provide a central motivation for this paper. Our results show that groups hyperbolic relative to such subgroups still satisfy the coarse Baum–Connes conjecture.
@@ 2 The twisted coarse Baum–Connes conjecture with respect to stable coarse algebras
In this section, we review some background on the coarse Baum–Connes conjecture together with its variants. More precisely, we will focus on a version of twisted coarse Baum-Connes conjecture with respect to certain stable coarse algebras. This is a new framework to study the coarse geometry of metric spaces with bounded geometry. Compared with the usual coarse Baum-Connes conjecture, the new framework captures more information of the large scale geometry of the considered space.
@@ 2.1 Twisted Roe algebras
    For each $x\in X$, we have that {f(x):f∈Γ(X, AK)}=AK;
@@
    For any two bounded sequences $a,b\in\ell^{\infty}(X,B(\mathcal{H}))$, and $f\in\ell^{\infty}(X;\mathcal{A}\otimes\mathcal{K})$, we define a new function $f^{\prime}\in\ell^{\infty}(X;\mathcal{A}\otimes\mathcal{K})$ by f’(x)=(id_A⊗a(x))f(x)(id_A⊗b(x)). Then $f\in\Gamma(X,\mathcal{A}\otimes\mathcal{K})$ implies $f^{\prime}\in\Gamma(X,\mathcal{A}\otimes\mathcal{K})$ for any two bounded sequences $a,b\in\ell^{\infty}(X,B(\mathcal{H}))$;
@@
    The propagation of $T$ is defined to be Propagation(T):=sup{d(x,y):T_x,y≠0}∈[0,∞]. We say that $T$ has finite propagation if ${\rm Propagation}(T)<\infty$.
@@
Next, we recall the definition of twisted Roe algebras. This notion was first introduced by the first-named author together with L. Guo in [DG24], where the coefficient algebras were taken to be coarse algebras. In the present paper, we instead use stable coarse algebras to define twisted Roe algebras. With this modification, the resulting twisted Roe algebras are invariant under coarse equivalences in a certain sense (Theorem 2.14).
We fix a net $X$ of $M$ and a Borel covering $\{B_{x}\}_{x\in X}$ satisfying the four conditions in Lemma 2.4. For every $x\in X$, fix a unitary
$U_{x}:\ell^{2}(B_{x}\cap M_{0})\otimes\mathcal{H}\to\mathcal{H}$ (1)
between Hilbert spaces. This induces an isomorphism on the Hilbert $A$-modules $U_{x}\otimes\text{id}_{\mathcal{A}}:\ell^{2}(B_{x}\cap M_{0})\otimes\mathcal{H}_{\mathcal{A}}\to\mathcal{H}_{\mathcal{A}}$.
For every $T=\left(T_{xy}\right)_{x,y\in M_{0}}\in B(\ell^{2}(M_{0},\mathcal{H}_{\mathcal{A}}))$ and every partial translation $v:D\to R$ on $X$, we define a bounded function
@@
by
@@
It will be shown in Lemma 2.13 that the definition of the $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$-twisted Roe algebras are independent of the choice of a Borel cover $\{B_{x}\}$ satisfying the four conditions in Lemma 2.4.
@@
    For the stable coarse $X$-algebra $\ell^{\infty}(X,\mathcal{A}\otimes\mathcal{K})$, we have that C^*(X, ℓ^∞(X, AK))=C^*(X,A).
@@
    For the stable coarse $X$-algebra $C_{0}(X,\mathcal{A}\otimes\mathcal{K})$, we have that C^*(X, C_0(X, K))≅AK.
@@
$f^{*}(\Gamma(Y,\mathcal{A}\otimes\mathcal{K}))=\{\varphi\circ f:\varphi\in\Gamma(Y,\mathcal{A}\otimes\mathcal{K})\}.$
Then in this case, the proof is the same as that of coarse invariance Roe algebras, and it is not necessary to require that the coarse algebras are stable. In Theorem 2.14, we prove a more general coarse invariance.
@@
It is obvious that $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$ is a $C^{*}$-subalgebra of $\ell^{\infty}(X,\mathcal{A}\otimes\mathcal{K})$. For any partial translation $v$ on $X$ and any $\varphi\in\Gamma(X,\mathcal{A}\otimes\mathcal{K})$, we show that $v^{*}\varphi\in\Gamma(X,\mathcal{A}\otimes\mathcal{K})$. For any subset $\tilde{Y}\subset f(X)$, we fix a set of representatives
@@
$(v^{*}\varphi)_{\tilde{X}}(y)=(v^{*}\varphi)(x_{y})=\varphi(v(x_{y}))$
and $(v^{*}\varphi)_{\tilde{X}}(y)=0$ if $y\notin\tilde{Y}$ or $x_{y}\notin\operatorname{Dom}(v)$.
@@
$v_{j}^{*}(\varphi_{\tilde{X}_{j}})(y)=\varphi_{\tilde{X}_{j}}(v_{j}(y))=\varphi_{\tilde{X}_{j}}(f(v(x_{y})))=\varphi(v(x_{y}))$
and $v_{j}^{*}(\varphi_{\tilde{X}_{j}})(y)=0$ if $y\notin A_{j}$. Therefore $v^{*}\varphi=\sum_{j=1}^{N}v_{j}^{*}(\varphi_{\tilde{X}_{j}})\in\Gamma(X,\mathcal{A}\otimes\mathcal{K})$. It is easy to see $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$ is stable if $\Gamma(Y,\mathcal{A}\otimes\mathcal{K})$ is. ∎
@@
Let $M$ be a metric space with bounded geometry, and let $X$ and $Y$ be nets of $M$ with associated Borel covers $\{B_{x}\}_{x\in X}$ and $\{C_{y}\}_{y\in Y}$ of $M$ satisfying the four conditions in Lemma 2.4. Suppose that $\Gamma(Y,\mathcal{A}\otimes\mathcal{K})$ is a stable coarse $Y$-algebra and $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$ is a coarse $X$-algebra induced by any coarse equivalence between $X$ and $Y$, which is close to the identity map on $M$, by the formula (4). Denote by $C^{*}(M;\Gamma(X,\mathcal{A}\otimes\mathcal{K}))$ (resp. $C^{*}(M;\Gamma(Y,\mathcal{A}\otimes\mathcal{K}))$) the $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$-twisted (resp. $\Gamma(Y,\mathcal{A}\otimes\mathcal{K})$-twisted) Roe algebra associated to the Borel cover $\{B_{x}\}_{x\in X}$ (resp. $\{C_{y}\}_{y\in Y}$). Then we have an equality
@@
Let $M$ and $N$ be metric spaces with bounded geometry, and let $X\subset M$ and $Y\subset N$ be nets of $M$ and $N$, respectively. Let $f:M\to N$ be a coarse equivalence, and fix Borel covers $\{B_{x}\}_{x\in X}$ and $\{C_{y}\}_{y\in Y}$ satisfying four conditions in Lemma 2.4, together with a coarse equivalence $g:X\to Y$ that is close to $f|_{X}$.
@@ 2.2 The twisted coarse Baum–Connes conjecture
Each Rips complex $P_{d}(X)$ can be equipped with the normalized spherical metric as follows. The simplex spanned by $\{x_{1},x_{2},\cdots,x_{n}\}$ is identified with the quadrant with diameter $1$:
$S^{n-1}_{+}=\left\{(a_{1},a_{2},\cdots,a_{n}):\sqrt{\sum_{i=1}^{n}a_{i}^{2}}=\frac{2}{\pi},~a_{i}\geq 0,\forall~1\leq i\leq n\right\}$
via the map
$\sum_{j=1}^{n}a_{j}x_{j}\mapsto\frac{2}{\pi}\left(\frac{a_{1}}{\sum_{i=1}^{n}a_{i}^{2}},\cdots,\frac{a_{n}}{\sum_{i=1}^{n}a_{i}^{2}}\right).$
Next, the Rips complex $P_{s}(X)$ is equipped with the path metric, and the distance between different connected components is defined to be infinity. For each $s>0$, we consider a countable dense subset $P_{r}(X)_{0}\subset P_{r}(X)$ such that $P_{s}(X)_{0}\subset P_{s^{\prime}}(X)_{0}$ if $s<s^{\prime}$.
@@
Since the right hand side of (6) is independent of $s$ by Theorem 2.14, passing to the limit yields the twisted coarse assembly map
@@
We first establish how a continuous and coarsely equivalent map $f:M\to N$ induces a map on the $K$-theory of their twisted localization algebras. This construction allows us to show the coarse invariance of the twisted coarse Baum–Connes conjecture. We proceed by recalling the necessary terminology (cf. [WY20, Definition 4.3.3]).
@@
Let $M$ and $N$ be metric spaces with bounded geometry. Fix their net $X\subset M$ and $Y\subset N$, and Borel covers $\{B_{x}\}_{x\in X}$ and $\{C_{y}\}_{y\in Y}$ of $M$ and $N$ satisfying the four conditions in Lemma 2.4, respectively. Let $f:M\to N$ be a continuous coarsely equivariant map. Fix any stable coarse $Y$-algebra $\Gamma(Y,\mathcal{A}\otimes\mathcal{K})$, and let $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$ be the coarse $X$-algebra induced by $f$.
We can also take nets $X^{\prime}\subset M$ and $Y^{\prime}\subset N$ and Borel covers $\{B^{\prime}_{x}\}_{x\in X^{\prime}}$ and $\{C^{\prime}_{y}\}_{y\in Y^{\prime}}$ of $M$ and $N$ such that $f$ is a bijection between $X^{\prime}$ and $Y^{\prime}$ and $f(B^{\prime}_{x})\subset C^{\prime}_{f(x)}$. We fix a decreasing sequence $\{\varepsilon_{j}\}_{j}$ of positive numbers with $\lim_{j}\varepsilon_{j}=0$. For each $x\in X^{\prime}$, since $f$ is continuous, by decomposing $C^{\prime}_{f(x)}$ into a smaller disjoint Borel sets satisfying the condition in Lemma 2.4, we obtain a unitary
@@
such that $\operatorname{supp}V_{j}(x)\subset\{(p,q)\in B^{\prime}_{x}\times C^{\prime}_{f(x)}:d(f(p),q)\leq\varepsilon_{j}\}$. We consider
@@
then $V_{j}$ is an $\varepsilon_{j}$-cover of $f$. We denote by $\Gamma(X^{\prime},\mathcal{A}\otimes\mathcal{K})$ and $\Gamma(Y^{\prime},\mathcal{A}\otimes\mathcal{K})$ the coarse algebra induced by the coarse equivalence between $X$ and $X^{\prime}$, and between $Y$ and $Y^{\prime}$, respectively. Then, it is obvious that the adjoint by $V_{j}$ sends $C^{*}(M;\Gamma(X^{\prime},\mathcal{A}\otimes\mathcal{K}))$ to $C^{*}(N;\Gamma(Y^{\prime},\mathcal{A}\otimes\mathcal{K}))$. By Lemma 2.13, for the originally fixed $\{B_{x}\}_{x\in X}$ and $\{C_{y}\}_{y\in Y}$, the adjoint by $V_{j}$ sends $C^{*}(M;\Gamma(X,\mathcal{A}\otimes\mathcal{K}))$ to $C^{*}(N;\Gamma(Y,\mathcal{A}\otimes\mathcal{K}))$. Define a uniformly continuous family of unitaries $(W(t))_{t\in[1,\infty)}$, where
@@
Then, $(W(t))_{t}$ is a continuous cover of $f$ and the adjoint by $(W(t))_{t}$ intertwines $C^{*}_{L}(M;\Gamma(X,\mathcal{A}\otimes\mathcal{K}))$ and $C^{*}_{L}(N;\Gamma(Y,\mathcal{A}\otimes\mathcal{K}))$. We denote
@@
$\text{ad}(W_{t})\left(C^{*}_{L}(M;\Gamma(X,\mathcal{A}\otimes\mathcal{K}))\right)\subset C^{*}_{L}(N;\Gamma(Y,\mathcal{A}\otimes\mathcal{K})).$